# A regularized continuum model for traveling waves and dispersive shocks of the granular chain

### v1

@@ 7.3 Initial data for the density variable $\rho$
Now, we fix the value of $\rho^{+}$ (for which the nature of the equations allows the freedom to arbitrarily select), and based on the jump condition of Eq. (7.9), we take the following initial condition for the variable $\rho$,
@@ 7.4 Initial data for the discrete model
To set up the Riemann problem for the original granular DDE that is consistent with the initial data we applied to the continuum model (See section 7.3 for details), we need to further understand an alternative first-order form of our continuum PDE which has a direct connection to the lattice system. In particular, we observe that instead of rewriting the continuum model as Eq. (7.4), it can also be written as follows,
@@
Here, $r(X,T)$ still represents an approximation of the strain of the lattice $r_{n}$, but $v(X,T)$ is related to the velocity of the displacement, $\dot{u}_{n}$. So both variables of the PDE formulation are directly related to physical variables of the granular chain. We then notice that, by setting $\rho=v-\frac{\epsilon^{2}}{12}v_{XX}$, the system (7.12) is equivalent to (7.4). This gives an interpretation of $\rho$ in terms of the physical variables of the granular chain (namely in terms of the velocity of the displacement).
@@
$v\left(X,0\right)$ $=\mathcal{F}^{-1}\left[\frac{\mathcal{F}\left[g(X)\right]}{1+\frac{1}{12}{\epsilon^{2}K^{2}}}\right],$ (7.14b)
where $\mathcal{F}^{-1},\mathcal{F}$ refer to the inverse and usual Fourier operators, and $K$ the Fourier wave number.
@@
where $s_{n}$ is the velocity in terms of the strain. Similar to the initial conditions in (7.14), the Riemann initial conditions for the discrete system (7.15) read
@@ 7.5 DSW fitting
Next, we consider special reductions of Whitham’s modulation system that we have derived in Sec. 6. These reductions will be used to obtain valuable information about the dispersive shock waves, such as the leading and trailing edge speeds as well as the amplitude of the solution at the leading edge of the DSW and the wavenumber at its trailing edge. To this end, we apply the well established, and highly successful in the continuum limit (yet not much explored in the lattice setting) “DSW fitting” method of El [22]. One preliminary and important task is to compute the associated linear dispersion relation of the original continuum model (2.12). Leveraging our earlier analysis in Eq. (2.13), we recall that the linear dispersion relation reads:
@@
$\log\left|\frac{\overline{r}}{r^{+}}\right|$ $=\frac{2}{p-1}\left(\sqrt{1+\frac{1}{12}\epsilon^{2}K^{2}}-\frac{1}{2}\log\left|1+\frac{1}{2}\sqrt{4+\frac{\epsilon^{2}K^{2}}{3}}\right|+\frac{1}{2}\log\left|\frac{4\left(1-\frac{1}{2}\sqrt{4+\frac{\epsilon^{2}K^{2}}{3}}\right)}{\epsilon^{2}K^{2}}\right|+\frac{1}{2}\log\left(12+\epsilon^{2}K^{2}\right)-1\right)$ (7.20a)
$\log\left|\frac{\overline{r}}{r^{-}}\right|$ $=\frac{2}{p-1}\left(\sqrt{1-\frac{1}{12}\epsilon^{2}\widetilde{K}^{2}}-\frac{1}{2}\log\left|1+\frac{1}{2}\sqrt{4-\frac{\epsilon^{2}\widetilde{K}^{2}}{3}}\right|+\frac{1}{2}\log\left|12-\epsilon^{2}\widetilde{K}^{2}\right|+\frac{1}{2}\log\left|\frac{4\left(1-\frac{1}{2}\sqrt{4-\frac{\epsilon^{2}\widetilde{K}^{2}}{3}}\right)}{\epsilon^{2}\widetilde{K}^{2}}\right|-1\right).$ (7.20b)
@@
An important observation here is that in expression (7.20) and (7.21), there is only a single effective parameter $\epsilon K$. This implies that once we return to the original granular variables and compare the wavenumber prediction of the DSW fitting and the wavenumber corresponding to the lattice, the prediction once again will be independent of $\epsilon$, in light of the relationship $k=\epsilon K$. The same is true for the leading and trailing edge speeds, namely they are independent of $\epsilon$, much like the case of the solitary waves, whose predictions were also independent of $\epsilon$.
@@ 8 Numerical validation
We are now ready to validate the theoretical results obtained previously by numerically solving the Riemann problems and comparing the numerical results with the associated quantities derived from the DSW fitting method.
@@
Figure 4 shows the granular lattice subject to the Riemann initial data given by Eq. (7.16) with $r^{-}=1$ and $r^{+}=0.95$ (see the red dots). Only a zoom of the right-moving wave is shown, which is where the DSW appears. The solid blue lines are the numerical simulation of the regularized continuum PDE proposed herein, namely Eq. (2.12) with the consistent initial data given by Eqs. (7.2) and  (7.10). The lattice and PDE simulations agree very well. The solid black lines represent the approximations obtained from the DSW fitting method described previously. In particular for panel (a), the vertical lines represent the trailing and leading edges, namely, $60s^{-}$ and $60s^{+}$, respectively, where $t=60$ is the time chosen and $s^{-},s^{+}$ are the speeds from Eq. (7.21). The sloped lines are an approximation of the envelope, and were obtained by connecting the point $(60s^{-},r^{-})$ with $(60s^{+},a^{+}+r^{+})$ (the top line) and the line defined by the points $(60s^{-},r^{-})$ with $(60s^{+},r^{+})$, (the bottom line). Panel (b) is similar but the time $t=150$ is used in place of $t=60$. The lattice simulations are encompassed by the triangle enclosed by the solid lines demonstrating that the theoretical prediction of the DSW characteristics (based on the DSW fitting discussed above) agrees quite well with the numerical observations. Such a quantitatively adequate fitting (everywhere but for part of the DSW’s linear tail) also indicates that the proposed PDE model (2.12) succeeds in characterizing the DSW embedded in the discrete model. The inaccuracy at the tail is common when approximating DSWs [43]. While the linear tails decay in time (with the rate $t^{-1/3}$) [44] it will always be present in any finite simulation.
@@
$a^{u}$ $=r^{-}+\frac{\left|r^{-}-r^{+}\right|}{2},$ (8.2a)
$a^{l}$ $=r^{-}-\frac{\left|r^{-}-r^{+}\right|}{2}.$ (8.2b)
These values are approximately $3/4$ and $1/4$ the amplitude of the DSW, respectively. Then, the sets of local maxima and minima are those with values that fall in the intervals, respectively,
@@
Figure 5: The leading-edge quantities comparison: the left and the right panel refer to the comparison of the leading-edge speeds and soliton amplitude, respectively. The red solid line represents the theoretical prediction of the leading-edge quantities, while the blue dots and pink circles refer to the numerical leading-edge quantities obtained from the simulation of the regularlized continuum PDE and the lattice DDE, respectively. The background information $r^{-}=1$ is fixed, but $r^{+}$ is varied. Here $\epsilon=0.1$.
@@
We display first the comparison of the leading-edge speed in Fig. 5(a) for $r^{-}=1$ fixed and various $r^{+}$. The value $\epsilon=0.1$ is fixed (although recall that the prediction of the DSW fitting method is independent of $\epsilon$). The red line indicates the theoretical prediction from the DSW fitting method, see Eqs. (7.21), where the blue solid dots and magenta open circles are the measured speeds from the numerical solution of the PDE and DDE, respectively. This is done for three values of the nonlinearity $p=3/2,2,3$. In all three cases, the agreement is very good overall. As the jump height $|r^{+}-r^{-}|$ decreases, the agreement between the DSW fitting prediction and simulation becomes better. For all values of $r^{+}$ considered, the blue dots fall almost perfectly into the magenta circles, indicating the strong agreement between DDE and PDE simulations. Figure 5(b) is similar to panel (a), but now the leading-edge amplitude is shown. Once again the markers are from the DDE and PDE simulations. Here, we only show results for $p=2$ and $p=3$, since we do not have an explicit formula for the amplitude in the case $p=3/2$. The leading-edge amplitude formulas are from Eq. (4.9) for $p=2$ and Eq. (4.22) for $p=3$ where the speed $c$ is replaced by $s^{+}$ in the formulas. While the quantitative agreement is not as good for larger jump heights when compared to panel (a), the overall trends are similar. Considering both panels (a,b) together, we can see that as the jump height decreases, the DSW increases its speed, but its amplitude decreases. While this may seem initially counter-intuitive (based on the situation for solitary waves), it is actually quite natural, since the background of the leading edge solitary wave is increasing (as the jump height decreases).
Figure 6: The trailing-edge quantities comparison: the left and right panels depict the comparison of the trailing-edge speeds and wavenumbers, respectively. The red solid line refers to the theoretical prediction of the trailing-edge quantities, while the blue dots and the pink circles represent the numerical trailing-edge quantities computed from the simulation of the regularized PDE and lattice DDE, respectively. The background information $r^{-}=1$ is fixed, while the value of $r^{+}$ is varied.
@@
Next, we investigate the trailing-edge quantities comparison, shown in Fig. 6. In panel (a) the trailing edge speed is shown, where the formula used to obtain the trailing edge speed is from Eq. (7.21). We notice that the trailing edge speeds of the DSWs from the numerical simulation of the Riemann problems are obtained by estimating the slope of $x$-$t$ plane where $x$ denotes the coordinate of the trailing edge locations measured by the method described above, and $t$ the temporal coordinate which corresponds to the time that the numerical solution is saved. The same marker and line conventions are used as in the previous figure. First, we notice clearly that the agreement is not as favorable as the one we have seen for the leading edge. As discussed above, there is always the presence of a linear tail in the finite simulations, and this makes the estimation of the trailing edge more prone to estimate errors. Thus it is possible that the larger deviations between the numerical simulations and the DSW fitting predictions are due to the method used to estimate the trailing edge in the simulation, rather than to inaccuracies in the DSW fitting approach. Nevertheless, as we can observe from Figure 6(a), the theoretical and numerical computed quantities still agree quite reasonably in the sense that all of them are rather proximal and indeed do not deviate significantly from one another. Figure 6(b) is similar to panel (a), but now the wavenumber is shown. The formula used to obtain the trailing edge wavenumber is obtained by solving Eq. (7.20). On the other hand, to estimate the numerical trailing edge wavenumber in the simulation of the Riemann problems, we focus on the final snapshot of the numerical solution, and then find the two local peaks of the DSW which are right to the trailing edge location that is described previously. We measure the distance between these two adjencent peaks of the DSW as the wavelength denoted by $\lambda$, and finally we compute the numerical trailing edge wavenumber by the formula of $K=\frac{2\pi}{\lambda}$. Finally, it is also worthwhile to mention that we need to be careful about the fact that $K=\epsilon^{-1}k$ which shows how the wavenumbers of the continuum PDE (2.12) and the discrete granular model (2.2) are related. Hence, in the numerical simulation of the granular discrete system (2.2), we scale and use the computational domain that $X=\epsilon n$, where $n$ denotes the lattice site, so that the ultimate estimated wavenumber of the simulation of the discrete system is comparable to that of the continuum model.
@@
The overall agreement between the various wavenumbers is quite good, despite larger deviations, stemming, possibly, from the uncertainty in the numerical wavenumber estimation. Considering both panels (a,b) together, we can see that as the jump height decreases, the trailing edge speed increases (just like the leading edge speed) and the wavenumber decreases. It is interesting to note that, if we consider the wavenumber in terms of the granular variables (where the value of the wavenumber will be multiplied by $\epsilon=0.1$) all wavenumbers are below $\pi$. The latter would correspond to a binary oscillation in the lattice. While the wavenumbers are bounded in the discrete problem, namely, being confined to $[0,\pi]$, the wavenumber is unbounded in the PDE. Indeed, it has been conjectured in [17] (see also [45] which includes a rigorous treatment in the special case of the Toda lattice) that the wavenumber reaching $\pi$ within the core of the DSW represents a key change in the structure of the DSW. In that case, the DSW would connect to a periodic (binary) wave, rather than being connected to a constant. Such a feature could not be captured by the PDE model, and highlights the fact that larger errors are to be expected for larger wavenumbers and (according to Fig. 6(b)) larger jump heights.

### v2

@@ 7.2 Initial data for the continuum model
If considering step initial data, as defined in Eq. (7.8), one can freely select $r^{+},r^{-},\rho^{+}$ and then $\rho^{-}$ is determined via Eq. (7.7).
For numerical simulations, we will employ a spectral method to discretize the spatial variables, and thus we require initial conditions that will respect the periodic boundary conditions. The following “box-type” initial data are one such choice and are analogous to Eq. (7.8),
$r(X,0)=\begin{cases}r^{-}\,,&a<X<b\,,\\ r^{+}\,,&X<a~{}\vee~{}X>b\,,\end{cases}\qquad\rho(X,0)=\begin{cases}\rho^{-}\,,&a<X<b\,,\\ \rho^{+}\,,&X<a~{}\vee~{}X>b\end{cases}$ (7.8)
where $a,b\in\mathbb{R}$ are two real constants with $a<b$. The discontinuity in this initial data, however, makes it unsuitable for computations. Thus, for the numerical approximation of the PDEs, we employ a smooth approximation of the ”box-type” initial strain,
We now need to find an appropriate smooth approximation of $\rho(X,0)$ that is consistent with Eq. (7.8) and that satisfies the jump conditions Eq. (7.7). Assuming $r^{+},r^{-},\rho^{+}$ are fixed (for which the nature of the equations allows the freedom to arbitrarily select), and based on the jump condition of Eq. (7.7), we take the following initial condition for the variable $\rho$,
@@
With $r_{-}>r_{+}$ and $\rho(X,0)$ chosen according to Eq. (7.7), a right-moving DSW will form on the right of the ”box”, whereas a left-moving rarefaction wave will form on the left of the ”box”. Since our focus is on the formation of DSWs, we do not report results on the rarefaction waves. See Secs. 7.4 and 8 for examples of right-moving DSWs.
If one were to demand that the Riemann invariant $q_{1}$ be fixed (rather than $q_{2}$) the jump condition Eq. (7.7) would become
$\rho^{-}-\frac{2\sqrt{p}}{p+1}(r^{-})^{\frac{p+1}{2}}=\rho^{+}-\frac{2\sqrt{p}}{p+1}(r^{+})^{\frac{p+1}{2}}.$ (7.12)
It is worth noting that a choice of $r_{-}>r_{+}$ and $\rho(X,0)$ to satisfy this jump condition would lead to a left-moving DSW on the left of the ”box” and a rarefaction wave moving to the right. The results are qualitatively similar to the case with $q_{2}$ fixed, and so this setting is not discussed further herein.
@@ 7.3 Initial data for the discrete model
To set up the Riemann problem for the original granular DDE that is consistent with the initial data we just discussed we need to further understand how the density variable $\rho$ relates to the discrete variables. To this end, we consider an alternative first-order form of our continuum PDE which has a direct connection to the lattice system. In particular, we observe that instead of rewriting the continuum model as Eq. (7.1), it can also be written as follows,
@@
Here, $r(X,T)$ still represents an approximation of the strain of the lattice $r_{n}$, but $v(X,T)$ is related to the particle velocity, $\dot{u}_{n}$. So now both variables of the PDE formulation are directly related to physical variables of the granular chain. We then notice that, by setting $\rho=v-\frac{\epsilon^{2}}{12}v_{XX}$, the system (7.13) is equivalent to (7.1). This gives an interpretation of $\rho$ in terms of the physical variables of the granular chain (namely in terms of the particle velocity).
@@
[…]
where $\mathcal{F}^{-1},\mathcal{F}$ refer to the inverse and usual Fourier operators, $K$ the Fourier wave number, and
$h(X)=\mathcal{F}^{-1}\bigg{[}\frac{1}{1+\frac{1}{12}{\epsilon^{2}K^{2}}}\bigg{]}=(\sqrt{3}/\epsilon)\,e^{-2\sqrt{3}|X|/\epsilon}\,.$ (7.16)
Figure 3 showcases the profiles of the initial conditions in Eq. (7.15).
@@
where $s_{n}$ is the strain derivative. Similar to the initial conditions in (7.15), the Riemann initial conditions for the discrete system (7.17) read
@@
Figure 3: The initial conditions in Eqs. (7.15). The background and parameter values are $r^{-}=1$, $r^{+}=0.9$, $p=2$, and $\epsilon=0.1$.
7.4 Numerical simulation and DSWs
Figure 4 displays the simulation of the Riemann problem for the DDE. A waveform emerges at the right upper corner of the ”box-type” initial data. This waveform connects left state ($r^{-}$) to the right state ($r^{+}$) and expands as time increases. If one zooms into a small spatial-window of the waveform, it will appear as though the wave is periodic. The underlying parameters of the periodic wave will depend on where in the lattice you zoom, and hence the waveform is modulated. This is the so-called DSW. At the front of the wave (called the leading edge) the wave closely resembles a solitary wave, and it travels with a near constant speed $s^{+}$ and has some amplitude $a^{+}$. On the other hand, the back part of the DSW (called the trailing edge) is characterized by a linear wave with some wavenumber $k^{-}$ traveling at a speed $s^{-}$ with amplitude that vanishes as the trailing edge is approached from within the DSW. For any finite time snapshot of the DSW, there will be linear waves that are located around the background value of $r^{-}$. These linear waves are not a part of the DSW per se, and vanish for $t\rightarrow\infty$. To avoid any possible confusion, we point out that the PDE (2.12) is non-dispersive only for small-amplitude waves on a zero background. Conversely, it is fairly straightforward to see that the PDE does admit small-amplitude waves riding on top of any non-zero background, which is what we refer to with the term linear waves. Figure 4(a) shows a time snapshot of the DSW, where the key elements of the DSW, namely the leading and trailing edge can be seen, along with the linear waves that are always present for any finite time snapshot. Figure 4(b) shows a plot of the strain magnitude $r(X,T)$ where the expanding oscillatory region (i.e., the DSW) can be seen. The dashed lines in panel (b) are analytical approximations of the leading and trailing edges that stem from the so-called DSW fitting method. Before discussing those approximations, we discuss some preliminaries involving the the dispersionless system.
Figure 4: Numerical simulation of the Riemann problem: The left panel shows the profile of the DSW of the continuum model (2.12) at $T=50$, while the right panel depicts the plot of the strain magnitude $r(X,T)$, and the two black dashed lines in the density plot represent the leading (upper) and the trailing (lower) edges of the DSW, respectively. In particular, the upper and the lower dashed black lines depict $X=s^{+}T$ and $X=s^{-}T$ where $s^{+},s^{-}$ are obtained based on Eqs. (7.23). Meanwhile, notice that the parameters of $r^{+},r^{-}$ in the Riemann initial condition (7.8) are set to be $r^{+}=0.8$, and $r^{-}=1$ in the numerical simulation.
@@ 7.5 DSW fitting
Next, we consider special reductions of Whitham’s modulation system that we have derived in Sec. 6. These reductions will be used to obtain valuable information about the dispersive shock waves, such as the leading and trailing edge speeds as well as the amplitude of the solution at the leading edge of the DSW and the wavenumber at its trailing edge. To this end, we apply the well established, and highly successful in the continuum limit “DSW fitting” method of El [22], which has not been much explored in the lattice setting.
We note that the applicability of the method requires the underlying modulation equations to be strictly hyperbolic and genuinly nonlinear, e.g., see [22, 45, 46]. It is not known at present whether the modulation equations derived in Section 6 satisfy these requirements. On the other hand, we will proceed under the assumption that these conditions are satisfied, and we will validate the calculations a posteriori by comparing the theoretical predictions with the results of direct numerical simulations of both the PDE (2.12) and the original DDE (2.2), demonstrating good agreement.
One preliminary and important task is to compute the associated linear dispersion relation of the original continuum model (2.12). Leveraging our earlier analysis in Eq. (2.13), we recall that the linear dispersion relation reads:
@@
$\log\left|\frac{\overline{r}}{r^{+}}\right|=\frac{2}{p-1}$ $\left(\sqrt{1+\frac{1}{12}\epsilon^{2}K^{2}}-\frac{1}{2}\log\left|1+\frac{1}{2}\sqrt{4+\frac{\epsilon^{2}K^{2}}{3}}\right|+\right.$ (7.22)
$\left.\frac{1}{2}\log\left|\frac{4\left(1-\frac{1}{2}\sqrt{4+\frac{\epsilon^{2}K^{2}}{3}}\right)}{\epsilon^{2}K^{2}}\right|+\frac{1}{2}\log\left(12+\epsilon^{2}K^{2}\right)-1\right)$
$\log\left|\frac{\overline{r}}{r^{-}}\right|=\frac{2}{p-1}$ $\left(\sqrt{1-\frac{1}{12}\epsilon^{2}\widetilde{K}^{2}}-\frac{1}{2}\log\left|1+\frac{1}{2}\sqrt{4-\frac{\epsilon^{2}\widetilde{K}^{2}}{3}}\right|+\right.$
$\left.\frac{1}{2}\log\left|12-\epsilon^{2}\widetilde{K}^{2}\right|+\frac{1}{2}\log\left|\frac{4\left(1-\frac{1}{2}\sqrt{4-\frac{\epsilon^{2}\widetilde{K}^{2}}{3}}\right)}{\epsilon^{2}\widetilde{K}^{2}}\right|-1\right).$
@@
An important observation here is that in expression (7.22) and (7.23), there is only a single effective parameter $\epsilon K$. This implies that once we return to the original granular variables and compare the wavenumber prediction of the DSW fitting and the wavenumber corresponding to the lattice, the prediction once again will be independent of $\epsilon$, in light of the relationship $k=\epsilon K$. The same is true for the leading and trailing edge speeds, namely they are independent of $\epsilon$, much like the case of the solitary waves, whose predictions were also independent of $\epsilon$. Finally, we recall that we can also compute the analytical prediction for the soliton amplitude for the cases of $p=2,3$ by using the formulas in Eq. (4.9) and Eq. (4.22), respectively, with replacing $c$ by $s^{+}$ in Eqs. (7.23).
@@ 8 Numerical validation
We are now ready to validate the theoretical results obtained previously by numerically solving the Riemann problems and comparing the numerical results with the associated quantities derived from the DSW fitting method. We employed a fourth-order Runge-Kutta method for time stepping both the discrete model (2.2) and the quasi-continuum PDE (2.12), as well as pseudo-spectral methods to discretize the spatial derivatives in Eq. (2.12).
@@
Figure 5 shows the granular lattice subject to the Riemann initial data given by Eq. (7.18) with $r^{-}=1$ and $r^{+}=0.95$ (see the red dots). Only a zoom of the right-moving wave is shown, which is where the DSW appears. The solid blue lines are the numerical simulation of the regularized continuum PDE proposed herein, namely Eq. (2.12) with the consistent initial data given by Eqs. (7.9) and  (7.10). The lattice and PDE simulations agree very well. The solid black lines represent the approximations obtained from the DSW fitting method described previously. In particular for panel (a), the vertical lines represent the trailing and leading edges, namely, $60s^{-}$ and $60s^{+}$, respectively, where $t=60$ is the time chosen and $s^{-},s^{+}$ are the speeds from Eq. (7.23). The sloped lines are an approximation of the envelope, and were obtained by connecting the point $(60s^{-},r^{-})$ with $(60s^{+},a^{+}+r^{+})$ (the top line) and the line defined by the points $(60s^{-},r^{-})$ with $(60s^{+},r^{+})$, (the bottom line). Panel (b) is similar but the time $t=150$ is used in place of $t=60$. The lattice simulations are encompassed by the triangle enclosed by the solid lines demonstrating that the theoretical prediction of the DSW characteristics (based on the DSW fitting discussed above) agrees quite well with the numerical observations.
Such a quantitatively adequate fitting (everywhere but for part of the DSW’s linear tail) also indicates that the proposed PDE model (2.12) succeeds in characterizing the DSW embedded in the discrete model. The inaccuracy at the tail is common when approximating DSWs [39]. While the linear tails decay in time (with the rate $t^{-1/3}$) [47] it will always be present in any finite simulation.
@@
$a^{u}$ $=r^{-}+\frac{\left|r^{-}-r^{+}\right|}{N},$ (8.2a)
$a^{l}$ $=r^{-}-\frac{\left|r^{-}-r^{+}\right|}{N},$ (8.2b)
where $N\in\mathbb{Z}$ is a positive integer. These values are approximately $3/4$ and $1/4$ the amplitude of the DSW, respectively. Then, the sets of local maxima and minima are those with values that fall in the intervals, respectively,
@@
Figure 6: The leading-edge quantities comparison: the left and the right panel refer to the comparison of the leading-edge speeds and soliton amplitude, respectively. The red solid line represents the theoretical prediction of the leading-edge quantities, while the blue circles and green squares refer to the numerical leading-edge quantities obtained from the simulation of the regularlized continuum PDE and the lattice DDE, respectively. The background information $r^{-}=1$ is fixed, but $r^{+}$ is varied. Here $\epsilon=0.1$.
@@
We display first the comparison of the leading-edge speed in Fig. 6(a) for $r^{-}=1$ fixed and various $r^{+}$. The value $\epsilon=0.1$ is fixed (although recall that the prediction of the DSW fitting method is independent of $\epsilon$). The red line indicates the theoretical prediction from the DSW fitting method, see Eqs. (7.23), where the blue circles and green squares are the measured speeds from the numerical solution of the PDE and DDE, respectively. This is done for three values of the nonlinearity $p=3/2,2,3$. In all three cases, the agreement is very good overall. As the jump height $|r^{+}-r^{-}|$ decreases, the agreement between the DSW fitting prediction and simulation becomes better. For all values of $r^{+}$ considered, the blue dots fall almost perfectly into the green squares, indicating the strong agreement between DDE and PDE simulations. Figure 6(b) is similar to panel (a), but now the leading-edge amplitude is shown. Once again the markers are from the DDE and PDE simulations. Here, we only show results for $p=2$ and $p=3$, since we do not have an explicit formula for the amplitude in the case $p=3/2$. The leading-edge amplitude formulas are from Eq. (4.9) for $p=2$ and Eq. (4.22) for $p=3$ where the speed $c$ is replaced by $s^{+}$ in the formulas.
While the quantitative agreement is not as good for larger jump heights when compared to panel (a), the overall trends are similar. Considering both panels (a,b) together, we can see that as the jump height decreases, the DSW increases its speed, but its amplitude decreases. While this may seem initially counter-intuitive (based on the situation for solitary waves), it is actually quite natural, since the background of the leading edge solitary wave is increasing (as the jump height decreases).
Figure 7: The trailing-edge speeds comparison: Note that the panels (a), (b), and (c) refer to the leading-edge speeds comparison for the cases of $p=3/2,2,3$, respectively. In addition, the solid red curves depict the analytical prediction of the trailing-edge speeds based on the formula of $s^{-}$ in Eqs. (7.23), while the blue circles and green squares showcase the numerically measured trailing-edge speeds of the continuum PDE (2.12) and the associated DDE (2.2), respectively.
Figure 8: The trailing-edge wavenumbers comparison: Notice that the red solid curves, from the top to the bottom, depict the analytical prediction of the trailing-edge wavenumbers, while the blue circles and green squares refer to the numerically measured trailing-edge wavenumbers of the continuum PDE model (2.12) and the corresponding discrete granular model (2.2).
@@
Next, we investigate the trailing-edge quantities comparison, shown in Figs. 7-8. In Fig. 7 the trailing edge speed is shown, where the formula used to obtain the trailing edge speed is from Eq. (7.23). We notice that the trailing edge speeds of the DSWs from the numerical simulation of the Riemann problems are obtained by estimating the slope of $x$-$t$ plane where $x$ denotes the coordinate of the trailing edge locations measured by the method described above, and $t$ the temporal coordinate which corresponds to the time that the numerical solution is saved. The same marker and line conventions are used as in the previous figure. First, we notice clearly that the agreement is not as favorable as the one we have seen for the leading edge. As discussed above, there is always the presence of a linear tail in the finite simulations, and this makes the estimation of the trailing edge more prone to estimate errors. Thus it is possible that the larger deviations between the numerical simulations and the DSW fitting predictions are due to the method used to estimate the trailing edge in the simulation, rather than to inaccuracies in the DSW fitting approach. Nevertheless, as we can observe from both Figs. 7-8, the theoretical and numerical computed quantities still agree quite reasonably in the sense that all of them are rather proximal and indeed do not deviate significantly from one another. Moreover, no signatures of non-monotonicity, e.g., of the trailing edge speed on the jump height are observed here (a feature that has been observed elsewhere, such as, e.g., [48]). Figure 8 is similar to Fig. 7, but now the wavenumber is shown. The formula used to obtain the trailing edge wavenumber is obtained by solving Eq. (7.22). To compute the wavenumber in the simulations, we first compute the inverse of the phase speed by comparing the time series of two consecutive nodes near the trailing edge. The inverse phase speed is then multiplied by the frequency to give an approximation of the wavenumber.
Notice that we have also conducted the numerical experiment with the initial condition (7.9) with the number $50$ replaced by $10$ and we do not find this change has made an appreciable effect on the numerically measured trailing-edge wavenumber.
Finally, it is also worthwhile to mention that we need to be careful about the fact that $K=\epsilon^{-1}k$ which shows how the wavenumbers of the continuum PDE (2.12) and the discrete granular model (2.2) are related. Hence, in the numerical simulation of the granular discrete system (2.2), we scale and use the computational domain $X=\epsilon n$, where $n$ denotes the lattice site, so that the estimated wavenumber in the discrete system can be compared to that of the continuum model.
@@
The overall agreement between the various wavenumbers is quite good, despite larger deviations, stemming, possibly, from the uncertainty in the numerical wavenumber estimation. Considering both panels (a,b) together, we can see that as the jump height decreases, the trailing edge speed increases (just like the leading edge speed) and the wavenumber decreases. It is interesting to note that, if we consider the wavenumber in terms of the granular variables (where the value of the wavenumber will be multiplied by $\epsilon=0.1$) all wavenumbers are below $\pi$. The latter would correspond to a binary oscillation in the lattice. While the wavenumbers are bounded in the discrete problem, namely, being confined to $[0,\pi]$, the wavenumber is unbounded in the PDE. Indeed, it has been conjectured in [17] (see also [49] which includes a rigorous treatment in the special case of the Toda lattice) that the wavenumber reaching $\pi$ within the core of the DSW represents a key change in the structure of the DSW. In that case, the DSW would connect to a periodic (binary) wave, rather than being connected to a constant. Such a feature could not be captured by the PDE model, and highlights the fact that larger errors are to be expected for larger wavenumbers and (according to Fig. 8) larger jump heights.